OpenStudy (anonymous):

Differential equation Find the differential equation for the family of curve cy + lnx = 0

OpenStudy (zzr0ck3r):

? take the derivative of that?

OpenStudy (anonymous):

I think taking the derivative is one of the steps. But I don't really know how to do it (I mean solving the whole question)

OpenStudy (zzr0ck3r):

you cant solve that

OpenStudy (zzr0ck3r):

3 variables in 1 equation

OpenStudy (anonymous):

I meant I can't find the differential equation.

OpenStudy (zzr0ck3r):

paste exactly what the book says to do

OpenStudy (anonymous):

This is the exact question - no missing words.

OpenStudy (zzr0ck3r):

i dif eq has the form dy/dx =.... or something like that, but you must know with what it is respect to

OpenStudy (zzr0ck3r):

you must know what variable to take the derviative in respect to.

OpenStudy (anonymous):

It didn't mention in the book, but I think it is w.r.t. x

OpenStudy (zzr0ck3r):

did it say family curve or family of curves?

OpenStudy (anonymous):

Ouch! family of curve. Sorry for that - one word missing

OpenStudy (zzr0ck3r):

so y = -lnx/c

OpenStudy (zzr0ck3r):

so y = -1/c * lnx dy/dx = -1/c * d(lnx)/dy

OpenStudy (zzr0ck3r):

can you do that?

OpenStudy (anonymous):

For the right side, what should it be -1/c * d(lnx)/dy ?

OpenStudy (anonymous):

Aren't we diff. w.r.t x?

OpenStudy (zzr0ck3r):

dx/dy is in respect to y

OpenStudy (anonymous):

Yes. But why did you diff. the left sides with respect to x while diff the right side with respect to y?

OpenStudy (zzr0ck3r):

o sorry that should be dx

OpenStudy (anonymous):

So, \(\frac{dy}{dx} = -\frac{1}{c} \times \frac{d(lnx)}{dx}\) \(\frac{dy}{dx} = -\frac{1}{cx}\) And then?

OpenStudy (anonymous):